# Weaving the (AdS) spaces with partial entanglement entropy threads

### v3

@@ 6.1 Conclusion
In summary, based on the PEE structure of a CFT and its geometrization scheme using the bulk geodesics anchored on the boundary, we can obtain a network of geodesics (or PEE threads) with the density of geodesics specified everywhere in the bulk [1]. In this paper we showed that, if we cut the AdS space open along any co-dimension 2 surface $\Sigma$, the density of cuts of the bulk geodesics is always $1/4G$ at any point on $\Sigma$. In other words, the AdS is weaved by geodesics, with the geodesics distributed uniformly at any point and along any direction. Then the area of any co-dimension 2 surface in AdS can be reconstructed by counting the number of intersections between the surface and the PEE network. Our study indicates that, the bulk geodesics are the basic elements that build up the AdS space in AdS/CFT.
Based on the PEE network, we provide a complete reformulation of the RT formula, which identifies the RT surface of an boundary region $A$ as the surface $\Sigma_{A}$ homologous to $A$ and has the minimal number of intersections with the PEE network. Our reformulation of the RT formula is partially inspired by the computation of the entanglement entropy in tensor network models, which is given by minimal number of cuts between a homologous path and the tensor network. Here the PEE network is analogous to the tensor network, and the number of intersections $\mathcal{N}(\Sigma_{A})$ is akin to the number of cuts in the tensor network. The PEE network can be viewed as a tensor network that precisely captures the entanglement structure of the boundary CFT at large $c$ limit.
@@
It is worth mentioning that, there have already been extensive studies (see [22, 70, 71, 72, 73, 74] for examples) on exploring holographic geometry reconstruction via the Crofton formula. Especially in [22] it was shown that, in AdS3 the differential entropy formula can be derived from the Crofton formula, and any bulk geodesic indeed represents the conditional mutual information between the degrees of freedom on the two boundary endpoints. Since it has been shown in [47] that, the conditional mutual information in CFT2 is exactly the PEE computed by the ALC proposal, our results exactly reproduce those in [22]. Nevertheless, most of these analyses in the above literature are confined to a time slice in AdS3. Remarkably, our work gives a more complete formalism for the area reconstruction in holography based on the partial entanglement entropy and generalize the geometry reconstruction program to AdS space in general dimensions.
It is also worth mentioning that, in [61] the authors introduced the so-called extensive mutual information (EMI), which is indeed the PEE we have defined. The EMI is also defined to satisfy the set of physical requirements we used to define the PEE. In addition, the authors of [61] assumed that the EMI that solves these requirements should be a mutual information, which is the essential difference between the EMI and the PEE. Nevertheless, the PEE obtained by solving the physical requirements only coincide with the mutual information in two-dimensional free fermions. In more generic theories, the PEE $\mathcal{I}(A,B)$ is not the mutual information, hence the EMI does do exist [75, 40, 42].
@@
Although our discussion is confined to a static time slice of Poincaré AdS, it can reconstruct the area of the RT surfaces for a generic boundary region in general dimensions. Our configuration can also be used to reconstruct the area of a generic co-dimension 2 surface, which goes beyond the RT formula. Furthermore, we can extract more information from the PEE thread configurations than the entanglement entropy, for example, the mixed state correlation represented by the EWCS [54] and the entanglement contour[60].
@@ 6.2 Future directions
Space with a black hole: There are cases where not all of the geodesics anchor on the boundary. For example, let us consider a time slice of the BTZ black hole. The boundary is in a mixed state, which is consistent with the fact that there are PEE threads emanating from the boundary, passing through black hole horizon and never come back to the boundary. Such PEE threads represent the thermal entropy of the boundary state. One may consider a canonical purification of the boundary which is dual to the eternal black hole, and study the configuration of the bulk PEE threads.
@@
PEE network as a tensor network toy model of quantum gravity: Compared to previous tensor network toy models of gravity, the PEE network is a well-defined continuous network that naturally extends to higher dimensions. It will be interesting to add bulk degrees of freedom to the PEE network to study quantum corrections [9] and quantum error correction properties [79, 32, 80] of the PEE network. Perhaps more property of quantum gravity could be studied based on the PEE network.
Acknowledgements
Y. Lu is supported by the National Natural Science Foundation of China under Grant No.12247161, the NSFC Research Fund for International Scientists (Grant No. 12250410250), China Postdoctoral Science Foundation under Grant No.2022TQ0140, and Shanghai Institute for Mathematics and Interdisciplinary Sciences. J. Lin is supported by the National Natural Science Foundation of China under Grant No.12247117. Q. Wen would like to thank Bartlomiej Czech, Veronika E. Hubeny, Ling-Yan Hung, Huajia Wang and Zhenbin Yang for helpful discussions.

### v4

@@ 6.1 Conclusion
In summary, based on the PEE structure of a CFT and its geometrization scheme using the bulk geodesics anchored on the boundary, we can obtain a network of geodesics (or PEE threads) with the density of geodesics specified everywhere in the bulk [66]. In this paper we showed that, if we cut the AdS space open along any co-dimension 2 surface $\Sigma$, the density of cuts of the bulk geodesics is always $1/4G$ at any point on $\Sigma$. In other words, the AdS is woven by geodesics, with the geodesics distributed uniformly at any point and along any direction. Then the area of any co-dimension 2 surface in AdS can be reconstructed by counting the number of intersections between the surface and the PEE network.
Based on the PEE network, we provide a reformulation of the RT formula, which identifies the RT surface of an boundary region $A$ as the surface $\Sigma_{A}$ homologous to $A$ and has the minimal number of intersections with the PEE network. Our reformulation of the RT formula is partially inspired by the computation of the entanglement entropy in tensor network models, which is given by minimal number of cuts between a homologous path and the tensor network. Here the PEE network is analogous to the tensor network, and the number of intersections $\mathcal{N}(\Sigma_{A})$ is akin to the number of cuts in the tensor network. The PEE network can be viewed as a tensor network that precisely captures the entanglement structure of the boundary CFT at large $c$ limit.
@@
6.2 Relation to other research drections
The extensive mutual information:
It is worth mentioning that, in [21] the authors introduced the so-called extensive mutual information (EMI), which is indeed the PEE we have defined if we do not require this quantity to be a mutual information [45]. The EMI is also defined to satisfy the set of physical requirements we used to define the PEE. In addition, the authors of [21] assumed that the EMI that solves these requirements should be a mutual information, which is the essential difference between the EMI and the PEE. Nevertheless, the PEE obtained by solving the physical requirements only coincide with the mutual information in two-dimensional free fermions. In more generic theories, the PEE $\mathcal{I}(A,B)$ is not the mutual information, hence the EMI does not exist [2, 92, 45].
@@ 6.2 Relation to other research drections
Differential entropy and Kinematic space
Although the computations and analyses in this paper may appear simple, we stress that the main results, i.e. the statement 2 and the reformulation of the RT formula for generic boundary regions, are a significant progress compared with our previous paper [66]. On the other hand, our analysis gives another derivation of the Crofton formula in AdS from a purely physical motivation. It is also worth mentioning that, there have already been extensive studies (see [27, 54, 95, 29, 58, 59, 55, 56, 57] for examples) on exploring holographic geometry reconstruction via the Crofton formula. Especially in [27] it was shown that, in AdS3 the differential entropy formula can be derived from the Crofton formula, and any bulk geodesic indeed represents the conditional mutual information between the degrees of freedom on the two boundary endpoints  [^8] Since it has been shown in [78] that, the conditional mutual information in CFT2 is exactly the PEE computed by the ALC proposal, our results exactly reproduce those in [27]. . Nevertheless, most of these analyses in the above literature are confined to a time slice in AdS3. It was unclear how to represent these geodesic using entropy quantities in higher dimensions, and here we give a satisfactory answer to this question: all the geodesics in the Crofton formula are represented by the two-point PEEs on the boundary CFT. Our work (at least) gives another formalism, or perhaps more natural formalism by generalizing the story to higher dimensions in a natural way, for the area reconstruction in holography based on the partial entanglement entropy.
Although our discussion is confined to a static time slice of Poincaré AdS, it can reconstruct the area of the RT surfaces for a generic boundary region in general dimensions. Our configuration can also be used to reconstruct the area of a generic co-dimension 2 surface, which goes beyond the RT formula. Furthermore, we can extract more information from the PEE thread configurations than the entanglement entropy, for example, the mixed state correlation represented by the EWCS [93] and the entanglement contour[24]. From the above discussions, we highlight three fundamental distinctions between our framework and kinematic space approach [6, 53, 25, 30, 27, 28]:
  • Conceptual foundations: The studies on Kinematic space are based on the Crofton formula in integral geometry, whereas our approach in some sense re-derived the Crofton formula in Poincare AdS. The two-point PEEs on the boundary are natural physical interpretation for the geodesics in Crofton formula.
  • Network interpretations: In the previous papers on Kinematic space, the discussion is mainly focused in the Kinematic space of the geodesics, while our discussion is mainly focused in the AdS space. In [28], the MERA tensor network was argued as a discretized embedding in the kinematic space. While in our construction, we take the PEE network as a continuous tensor network that make up the Poincare AdS space.
  • Higher dimension generalizations: In the previous papers on Kinematic space, while a geodesic in Poincare AdS3 space could be interpreted as a two-point conditional mutual information (CMI), the higher-dimensional generalizations are missing. Our framework resolves this by establishing a correspondence: the geodesics in general dimensional AdS space could be interpreted as the two-point PEEs, which coincide with two-point CMI in Poincare AdS3.
@@ 6.3 Future directions
It is interesting to observe that, the PEE threads emanating from the boundary fully foliate the Poincaré AdS space. But we think this is not an coincidence. The underlying reason for this complete foliation may be that the boundary state is pure, allowing the bulk geometry to be fully reconstructed from boundary data. Moreover, all threads must return to the boundary, as there should be no open endpoints in the bulk. If we extend the analysis to a BTZ black hole, threads originating from the boundary can only fully foliate the region outside the horizon (see forthcoming work [11]). In this case, some threads will enter the horizon, terminating at the singularity or extending to the second boundary in an eternal black hole setup. This reflects the thermal nature of the boundary state, with threads reaching the horizon contributing to the thermal entropy.
@@
where $\text{Geod}_{n}(M)$ is the space of all $n$-dimensional totally geodesic submanifolds $\Gamma$ and $c_{n,d}$ is some constant depending on the dimension of manifold $M$ and the dimension of the codimension-$n$ object $N$. The existence of totally-geodesic submanifolds of dimension $n\geqslant 2$ in a general Riemannian manifold is exceptional. But with the help of symmetries, we can find the measure $\mathrm{d}\Gamma$ and calculate the volume of $N$ with the Crofton formula, see [31] for an example in AdS3.
PEE network as a tensor network toy model of quantum gravity: Compared to previous tensor network toy models of gravity, the PEE network is a well-defined continuous network that naturally extends to higher dimensions. It will be interesting to add bulk degrees of freedom to the PEE network to study quantum corrections [39] and quantum error correction properties [5, 76, 46] of the PEE network. We refer to [87] for a recent development in this direction. The authors propose three new PEE tensor network models: namely, a factorized PEE, a Happy-like, and a random PEE tensor network. Furthermore, they prove that the number of intersection $\mathcal{N}_{\mathcal{E}_{A}}$ not only reproduce the area of the RT surface, but also yields the entanglement entropy for any boundary region.
Bulk metric reconstruction: In our framework, the bulk metric is given a prior, so that we can define a measure for PEE threads. However, one of the important problems in AdS/CFT is to derive bulk metric purely from the boundary CFT data. See [30, 8, 9, 62, 61] for relevant studies. It is possible that, matter fields or singularities can act as sources or sinks for PEE threads. Conversely, introducing such sources while still requiring a complete foliation could constrain the background metric in a manner consistent with Einstein’s equations, hence shed light on the bulk metric reconstruction program. We leave this for future investigations.
Acknowledgements
Y. Lu is supported by Shanghai Institute for Mathematics and Interdisciplinary Sciences. J. Lin is supported by the National Natural Science Foundation of China under Grant No.12505063. Q. Wen is supported by NSFC Grant No. 12447108. Q. Wen would like to thank Bartlomiej Czech, Veronika E. Hubeny, Ling-Yan Hung, Huajia Wang and Zhenbin Yang for helpful discussions.
Appendix A The Crofton formula
where $\overline{\text{Gr}}_{1}(\mathbb{R}^{2})$ is the space of geodesics in $\mathbb{R}^{2}$, $\mathrm{d}\Gamma$ is an appropriately normalized measure invariant under rigid motions (translations and rotations), and $(C\cap\Gamma)$ is the number of times a geodesic $\Gamma$ intersects with the curve $C$. The geodesics in $\mathbb{R}^{2}$ can be parameterized by
$x\cos\phi+y\sin\phi-p=0,$ (43)
$\mathrm{d}\Gamma=\mathrm{d}p\wedge\mathrm{d}\phi.$ (44)
It is easy to check this measure is invariant under rigid motions.
Assuming that $C$ is parameterized by $x=x(s),y=y(s)$, where the parameter $s$ is the length parameter. Let $\alpha$ denote the angle between the tangent at the point on $C$ and the straight line $\Gamma$ intersect $C$ at this point and $\theta$ the angle between the tangent to $C$ and $x$-axis. The measure expressed with parameters $s$ and $\alpha$ is then
$\mathrm{d}\Gamma=|\sin\alpha|\mathrm{d}s\wedge\mathrm{d}\alpha,$ (45)
The proof of the Crofton formula is now obvious
The Crofton formula can be generalized to Riemannian manifolds. Considering a $d$-dimensional Riemannian manifold $M$ (with a metric $g_{\mu\nu}$) and an arbitrary (d1)(d-1)-dimensional immersed hypersurface $\Sigma\subset M$, the Crofton formula states
where $\text{Geod}(M)$ is the space of all geodesics in $M$, $c$ is a constant and $\mathrm{d}\Gamma$ is a properly chosen measure on $\text{Geod}(M)$, which has to be invariant under changes of local coordinates and invariant under translations along geodesics.
@@ Appendix A The Crofton formula
The derivation of the Crofton formula in general Riemannian manifolds follows the similar strategy to that in the $\mathbb{R}^{2}$ case. We decompose the measure $\mathrm{d}\Gamma$ into two parts, one of which represents the position of all the points on the hypersurface $\Sigma$ and the other represents all the tangent direction of the geodesics passing through any point on $\Sigma$.
$\mathcal{L}=\sqrt{g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}},$ (48)
The co-sphere bundle $S^{*}M$ (i.e. pair $(x^{\mu},p_{\mu})$ with $p^{2}=1$) forms a manifold with $\dim=2d-1$, which is the phase space of $M$. And $(x^{\mu},p_{\mu})$ serves as a local coordinate system for $S^{*}M$, called the canonical coordinates. The geodesics $\Gamma$ of M{M} define a foliation $F_{\Gamma}$ of $S^{*}M$, in which the leaves $S^{*}M/F_{\Gamma}$ with dimension 2d22d-2. The canonical two-form
$\mathrm{d}p_{\mu}\wedge\mathrm{d}x^{\mu}$ (50)
on $S^{*}M$ is invariant under local coordinate transformations and the displacements on the leaves of the foliation (i.e. independent of the parameter $\tau$) [81]. It follows that the following (2d2)(2d-2)-form on $S^{*}M$ is also invariant
$\mathrm{d}\Gamma$ $=(\mathrm{d}p_{\mu}\wedge\mathrm{d}x^{\mu})^{d-1}$ (51)
which defines an invariant measure of (2d2)(2d-2)-dimensional set of geodesics $\Gamma$. It is called kinematic measure. Also the space of all the geodesics in the manifold is called the kinematic space [30, 27, 28, 26].
In other words, the bulk geodesics are parameterized by the position where they intersect with $\Sigma$ and the intersection direction captured by the d1d-1 independent conjugate momentum. To perform the integral over $\mathrm{d}\Gamma$, note that the conjugate momentum can be parameterized by the angle $\alpha_{i}$ between the momentum and $x^{i}$ axis: $p_{i}=\sqrt{g_{ii}}\cos\alpha_{i}$. Up to the sign, the kinematic measure becomes
where we recognize the area element $\mathrm{d}\sigma$ of a (d1)(d-1)-dimensional hypersurface $\Sigma$:
and the area element on the unit (d1)(d-1)-dimensional sphere in the direction $p_{d}$
Hence we can write
and perform integral for the form $\mathrm{d}\Gamma$ over $\Sigma$ and the upper half of the (d1)(d-1)-dimensional unit sphere [^9] The integral is over half the unit sphere because the geodesics are unoriented.
$=\mathrm{Area}(\Sigma)\frac{\Omega_{d-2}}{d-1},$
where $\Omega_{d-2}=\frac{2\pi^{(d-1)}/2}{\Gamma((d-1)/2)}$ is the area of the (d2)(d-2)-dimensional sphere and in the last step we apply the following relation